Circular No. 7444

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cq IN UGC 10354 A. V. Filippenko and A. L. Coil, University of California at Berke-

ley, report that a CCD spectrum (range 320–1000 nm) of SN 2000cq (cf. IAUC 7442), obtained on June 27 UT with the Shane 3-m reflector at Lick Observatory, reveals a broad Hα emission line at recession velocity 9000 km/s, equal to that listed by NED for the host galaxy. The Hα absorption component in this type-II supernova is weak, however.

SUPERNOVA 2000cr IN NGC 5395 R. T. Chornock, A. V. Filippenko, A. L. Coil, and D. C. Leonard,

University of California at Berkeley, write: “Inspection of a fully calibrated CCD spectrum (range 420–690 nm) of SN 2000cr obtained on June 27 UT with the Shane 3-m reflector at Lick Observatory reveals that it does not seem to be a type-II supernova as suggested by Jha et al. (IAUC 7443). There is no clear broad Hα emission, but a weak absorption trough at rest wavelength 637 nm may correspond to Hα with an expansion velocity of 9000 km/s. An even broader absorption trough at rest wavelength 613 nm might be Si II with an expansion velocity of 11 000 km/s, yet it is difficult to understand physically how this could exceed the Hα velocity. Another absorption line is at rest wavelength 434 nm. Finally, there is a very broad (FWHM = 30 nm) absorption trough centered on 487 nm, with perhaps two distinct minima at 480 and 493 nm. Among the spectra in our existing database, the spectrum of SN 2000cr most closely resembles that of the type-Ic supernova 1987M about a week past maximum brightness (Filippenko 1997, ARAA 35, 309), though SN 2000cr is probably younger. We surmise that SN 2000cr is a very young type-Ic supernova.”

CI AQUILAE CCD photometry by D. Hanžl, Brno, Czech Republic (Hipparcos com-

parison star HIP 92546): May 7.031 UT, V = 8.85 ± 0.02, B − V = +0.69 ± 0.02, V − I = +1.79 ± 0.03; 17.007, 9.51 ± 0.03, +0.74 ± 0.02, +1.67 ± 0.02; June 20.952, 12.49 ± 0.03, +0.44 ± 0.06, +1.38 ± 0.02.

COMET C/1999 S4 (LINEAR) Visual m_{1} and coma-diameter estimates: June 18.00 UT, 8.4, 4′ (M.

Reszelski, Szamotuly, Poland, 20×60 binoculars); 23.93, 8.5, 4′ (A. Baransky, Kiev, Ukraine, 12×80 binoculars); 25.98, 7.8, 5′ (Reszelski).

2000 June 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green